OpenStudy (anonymous):

calculate the frequency,energy and wavelength of the radiation corresponding to the spectral line to lowest frequency in lyman series in the spectra of hydrogen atom .also calculate the frequency fir the corresponding line in the spectra of Li+2 ion.[Rh=1.09678*10^7m,c=3*10^8m/s,h=6.62*10^-24Js]

OpenStudy (anonymous):

@Preetha

OpenStudy (frostbite):

Use the Rydberg formula. \[\large \frac{ 1 }{ \lambda }=R _{h}\left( \frac{ 1 }{ n ^{2} }-\frac{ 1 }{ m ^{2} } \right)\]

OpenStudy (frostbite):

When you have the wavelength λ, you can use the following relations to find: The frequency \(f\): \[\large f=\frac{ c }{ \lambda }\] The energy \(E\): \[\large E=hf\] Remember the Lyman series is when \(n=1\) so the Rydberg formula to use is: \[\large \frac{ 1 }{ \lambda }=R _{h}\left( 1-\frac{ 1 }{ m ^{2} } \right)\]

OpenStudy (frostbite):

Combine it all to one and we see: \[\large E _{photon}=hcR _{h}\left( \frac{ 1 }{ n ^{2} }-\frac{ 1 }{ m ^{2} } \right)\] Where: \[\large hcR _{h} \approx 13,6 \text{ eV}\]